magnetic field outside two parallel wires carrying a reverse current with a distane of D?
What are their conditions?
1
Expert's answer
2021-04-30T11:23:27-0400

Magnetic field:

"\overrightarrow{B_1}=\frac{\mu i}{4\pi D}(.)"


"\overrightarrow B_2=\frac{\mu i}{4\pi D}(.)"


Hence, the net magnetic field,

"\overrightarrow{B}=\overrightarrow{B_1}+\overrightarrow{B_2}"


"=\frac{2 \mu i}{4\pi D}(.)"